OpenStudy (anonymous):

Is 2(x^2 +4x -336) factorable?

OpenStudy (anonymous):

more than it is now? no, not that i know of.

OpenStudy (anonymous):

if you can find the factors of -336 that would add up to 4, then yes.

OpenStudy (anonymous):

16 and 21 are the factor, I would say, no! see why?

OpenStudy (anonymous):

Well, it actually depends on if you count irrational factors or not

OpenStudy (anonymous):

you would not get the factors that add up to 4 in this case, if you have factors that add up to 5, understand?

OpenStudy (anonymous):

@randerson4 and @zimmah, do you understand why you can't factor this?

OpenStudy (anonymous):

you can factor if the factors add up to the number next to x, in this case they don't.

OpenStudy (anonymous):

you can, but it won't be rational. you can't find rational numbers that add up to +4 and form 336 multiplied

OpenStudy (anonymous):

-336 even

OpenStudy (anonymous):

you can factor to 21 and -16, but you would get 5x, not 4x See?

OpenStudy (anonymous):

@zimmah, what the question means is into integers I think. I you right though.....

OpenStudy (anonymous):

Thanks! I am doing distance = rate x time problem, and I first arrived at 2x^2 + 8x -672 so I factored a 2 out of all. I just realized my mistake it should have been I believe 14 x, I am checking now.

OpenStudy (anonymous):

Yes but, \[(2*\sqrt{85}-2)*(2*\sqrt{85}+2)=336 \] and adding them would give 4*SQRT(85)

OpenStudy (anonymous):

@zimmah, yes, truly you can factor almost a/t, but I don't think the question meant that type of answer.

OpenStudy (anonymous):

true, that's why i initially said there was no answer, although i just wanted to make sure i didn't rule out the option of irrational solutions.

OpenStudy (anonymous):

K, that makes sense, but interger-wise it is unfactorable...

Directrix (directrix):

@randerson4 Which math course are you currently taking?

Directrix (directrix):

This is the first question @randerson4 has asked. He is a newbie. So, please don't make his first experience with OpenStudy unpleasant and confusing by arguing about factoring. Why not just say that the posted expression is NOT factorable over the set of Integers and wait for @randerson4 to follow up with a question of his own if he has one. I don't see the point of hijacking randerson's thread.

OpenStudy (anonymous):

@Directrix, that's what I said, wasn't it?

OpenStudy (anonymous):

No hard feelings. I made a silly mistake all is well I have solved my problem and obtained all the correct answers. It should have been 2(x^2 +14x-1176) lol. I pulled a Sheldon Cooper, that is I was careless. I am glad I found you guys! My next class is analytical geometry and calculus, so I am sure I will be back!

Directrix (directrix):

Good to hear @randerson4 . We look forward to seeing you here again.